Gideon is 30 yrs. younger than father. His mother is 3 times as old as his son. Find Gideons age if his father is 2 yrs. older than his mother.
Answer by mananth(16946)

let father be x years
Gideon = x-30
..
Mother = 3(x-30)
...
x=3(x-30)+2
x=3x-90+2
3x-x=90-2
2x=88
x= 44 fathers age
Gideon = 44-30
=14 years